# Correlation-driven branch in doped excitonic insulators

### v1

@@ Acknowledgements.
We thank K. Aido, M. Akiyama, J. Han, Y. Inokuma, T. Kondo, K. Kuroki, R. Mizuno, M. Ochi, Y. Ōno, and S. Tsuchida for fruitful discussions. This work was supported by Grants-in-Aid for Scientific Research from JSPS, KAKENHI Grant No. JP24K06939, No. JP24H00191, and No. JP24K01333. R.U. was supported by the Program for Leading Graduate Schools: “Interactive Materials Science Cadet Program” and JST SPRING, Grant No. JPMJSP2138. This project was made possible by the DLR Quantum Computing Initiative and the Federal Ministry for Economic Affairs and Climate Action; qci.dlr.de/projects/ALQU. The DMRG and TEBD simulations were performed using the ITensor library [73].

### v2

@@ Acknowledgements.
We thank K. Aido, M. Akiyama, J. Han, Y. Inokuma, T. Kondo, K. Kuroki, R. Mizuno, M. Ochi, Y. Ōno, and S. Tsuchida for fruitful discussions. This work was supported by Grants-in-Aid for Scientific Research from JSPS, KAKENHI Grant No. JP24K06939, No. JP24H00191, and No. JP24K01333. R.U. was supported by the Program for Leading Graduate Schools: “Interactive Materials Science Cadet Program” and JST SPRING, Grant No. JPMJSP2138. This project was made possible by the DLR Quantum Computing Initiative and the Federal Ministry for Economic Affairs and Climate Action; qci.dlr.de/projects/ALQU. S.E. acknowledges the Q-Neko project, which has received funding from the European Union’s Horizon Europe research and innovation programme under Grant Agreement No. 101241875. This work was also performed for Council for Science, Technology and Innovation (CSTI), Cross-ministerial Strategic Innovation Promotion Program (SIP), “Promoting the application of advanced quantum technology platforms to social issues” (Funding agency: QST). The DMRG and TEBD simulations were performed using the ITensor library [73, 74].
Data Availability
The data that support the findings of this article are openly available [75].